OpenStudy (celinegirl):

Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Laura has the movie. c(x) = 2.00x + 0.50 c(x) = 2.00 + 0.50x c(x) = 2.50x c(x) = (2.00 + 0.50)x

OpenStudy (mathstudent55):

The cost function is of the form: c(x) = fixed fee + variable fee The flat fee does not depend on the number of days. It is simply a fixed fee. The variable fee is variable because it depends on the number of days. At $0.50 per night, you'd pay $0.50 for 1 night, $0.50 * 2 = $1.00 for 2 nights, $0.50 * 3 = $1.50 for 3 nights, and $0.50 * x = 0.50x for x nights. Now add the fixed fee to the variable fee. Which cost function do you get?

OpenStudy (celinegirl):

either C or D.. idk it seems to confuse me

OpenStudy (celinegirl):

@mathstudent55 ^^